OpenStudy (anonymous):

today is my last day please help! 2x/5x+6=2/x What does x equal? x = -1 and x = 6 x = -1 and x = -6 x = 1 and x = -5 x = 1 and x = 5

OpenStudy (anonymous):

@reemii Can you help me please?

OpenStudy (reemii):

transform this and try to get an equation of the form \(x=...\). start by simplifying the first fraction.

OpenStudy (anonymous):

That would leave 3x.

OpenStudy (anonymous):

3x+6*

OpenStudy (reemii):

\(\frac{2x}{5x}=\frac25\)

OpenStudy (anonymous):

is that a three or a 5?

OpenStudy (reemii):

it's 2 divided by 5. look, you get: \[\frac25+6 = \frac2x\\ \frac{2 + 6\times 5}5 = \frac2x\\ x = \cdots\]